# NuGraph2 with Context–Aware Inputs: Physics–Inspired Improvements in Semantic Segmentation

### v1

@@ 3 Discussion
Tables 1 and 2 summarizes the results of all three interventions and the baseline network. The inclusion of extra input features, designed with physics intuition, proved to be the most effective intervention for our application. These features encoded non-local and structural information, such as node degree, local geometric context in the wirevstimewire\ vs\ time plane, and differential measures that reflect local track continuity, which significantly improved semantic classification metrics across all classes. In particular, Michel electron recall and precision showed the largest gains due to their previously poor performance caused by class imbalance and latent space overlap. This result highlights the value of embedding physics-informed context directly into the model’s input representation, as it helps disentangle class boundaries in the learned latent space.
@@ 4 Conclusions
Overall, our findings suggest that targeting Michel electrons in isolation is ineffective; progress arises when the latent representation is sharpened for all classes, allowing the minority class to benefit indirectly. Enhancing the model’s understanding of global class relationships and input structure benefits all classes, including Michel’s. For the NuGraph2 application, our results show the importance of integrating domain knowledge at the representation level, rather than solely through auxiliary objectives or task-specific losses.
@@ 5 Acknowledgments
The authors would like to thank V Hewes for their assistance in coding NuGraph extensions and understanding the intricacies of the original network’s implementation, and Burt Holtzman and Maria Acosta for their management of computing resources used for running a large amount of the analysis included in this work. We acknowledge the MicroBooNE Collaboration for making publicly available the data sets [18, 17] employed in this work.

### v3

@@ 3 Discussion
Tables 1 and 2 summarizes the results of all three interventions and the baseline network. The inclusion of extra input features, designed with physics intuition, proved to be the most effective intervention for our application. These features encoded non-local and structural information, such as node degree, local geometric context in the timevswiretime\ vs\ wire plane, and differential measures that reflect local track continuity, which significantly improved semantic classification metrics across all classes. In particular, Michel electron recall and precision showed the largest gains due to their previously poor performance caused by class imbalance and latent space overlap. This result highlights the value of embedding physics-informed context directly into the model’s input representation, as it helps disentangle class boundaries in the learned latent space.
@@ 4 Conclusions
Overall, our findings suggest that targeting Michel electrons in isolation is ineffective; progress arises when the latent representation is sharpened for all classes, allowing the minority class to benefit indirectly. Enhancing the model’s understanding of global class relationships and input structure benefits all classes, including Michels. For the NuGraph2 application, our results show the importance of integrating domain knowledge at the representation level, rather than solely through auxiliary objectives or task-specific losses.
@@ 5 Acknowledgments
The authors would like to thank Burt Holtzman and Maria Acosta for their management of computing resources used for running a large amount of the analysis included in this work. We acknowledge the MicroBooNE Collaboration for making publicly available the datasets [abratenko_2022_7261921, abratenko_2023_8370883] employed in this work.